Required information

Greenwood Company manufactures two products—14,000 units of Product Y and 6,000 units of Product Z. The company uses a plantwide overhead rate based on direct labor-hours. It is considering implementing an activity-based costing (ABC) system that allocates all of its manufacturing overhead to four cost pools. The following additional information is available for the company as a whole and for Products Y and Z:

Activity Cost Pool Activity Measure Estimated Overhead Cost Expected
Activity
Machining Machine-hours $ 198,000 10,000 MHs
Machine setups Number of setups $ 86,400 180 setups
Production design Number of products $ 82,000 2 products
General factory Direct labor-hours $ 248,000 10,000 DLHs
Activity Measure Product Y Product Z
Machining 7,800 2,200
Number of setups 50 130
Number of products 1 1
Direct labor-hours 7,800 2,200

11. Using the plantwide overhead rate, what percentage of the total overhead cost is allocated to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your answers to 2 decimal places.)

12. Using the ABC system, what percentage of the Machining costs is assigned to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your answers to 2 decimal places.)

13. Using the ABC system, what percentage of Machine Setups cost is assigned to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your answers to 2 decimal places.)

14. Using the ABC system, what percentage of the Product Design cost is assigned to Product Y and Product Z? (Round your answers to 2 decimal places.)

15. Using the ABC system, what percentage of the General Factory cost is assigned to Product Y and Product Z? (Round your intermediate calculations and final answers to 2 decimal places.)

i know its only four questions, but this the last question i need answered. ty
Answer #1

11) Percentage of plantwide overhead rate

Product Y Product Z
Plantwide overhead rate 7800/10000 = 78% 22%

12) Percentage of machining

Product Y Product Z
Machine hour 7800/10000 = 78% 22%

13) Percentage of machine setup

Product Y Product Z
Machine setup 50/180 = 27.78% 130/180 = 72.22%

14) Percentage of product design

Product Y Product Z
Product design 1/2 = 50% 50%

15) Percentage of Direct labor hours

Product Y Product Z
Direct labor hours 7800/10000 = 78% 22%